Grade 11Physical Chemistry

T the density of an ideal gas a is 1.43 g/l at STP. Determine the density of a at 17°c and 700 torr ( mm).

khusboo
8 Years agoGrade 11

1 Answer

Rituraj Tiwari
5 Years ago

To determine the density of the gas at 17°C and 700 torr, we will use the ideal gas law and the density formula:

Given Data:
Density at STP (P = 1 atm, T = 273 K) = 1.43 g/L
Temperature at new conditions, T₂ = 17°C = (17 + 273) K = 290 K
Pressure at new conditions, P₂ = 700 torr
Standard pressure, P₁ = 1 atm = 760 torr
Standard temperature, T₁ = 273 K
Step 1: Use the Density Formula for an Ideal Gas
The density of a gas is given by:

d = (P * M) / (R * T)

Since molar mass (M) and the universal gas constant (R) remain the same, the ratio of densities at different conditions can be written as:

(d₂ / d₁) = (P₂ * T₁) / (P₁ * T₂)

Step 2: Substitute the Given Values
d₂ / 1.43 = (700 * 273) / (760 * 290)

d₂ = 1.43 * (700 * 273) / (760 * 290)

Step 3: Solve for d₂
d₂ = 1.43 * (191100) / (220400)

d₂ = 1.43 * 0.8672

d₂ ≈ 1.24 g/L

Final Answer:
The density of the gas at 17°C and 700 torr is 1.24 g/L.